OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A. 2 B.- 4/5 C. 5/4 D.-2

OpenStudy (mayankdevnani):

If two points are given,\(\large \bf (x_1,y_1)~and~(x_2,y_2)\),then to find the SLOPE of line passing through them is ;- \[\large \bf Slope=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (mayankdevnani):

from this,can you solve your question ? @Archer3545

OpenStudy (anonymous):

not really

OpenStudy (mayankdevnani):

plug in the values in the equation shown above

OpenStudy (linn99123):

(1,-5)(4,1) (x1,y1)(x2,y2)

OpenStudy (mayankdevnani):

\[\large \bf Slope=\frac{1-(-5)}{4-1}=\frac{1+5}{3}=\frac{6}{3}=3\] Here :- \[\large \bf (x_1,y_1)=(1,-5)\] \[\large \bf (x_2,y_2)=(4,1)\]

OpenStudy (mayankdevnani):

here is the solution !!! @Archer3545

OpenStudy (mayankdevnani):

hope you understand, @Archer3545

OpenStudy (mayankdevnani):

sorry :) answer is 2

OpenStudy (mayankdevnani):

`you can reply here `

OpenStudy (mayankdevnani):